OpenStudy (anonymous):

The circumference of a circle is 31.4 feet. What is the area of the circle? Explain how you found your answer.

OpenStudy (anonymous):

First off you will need to know the formula for the area of a circle, which is A=pi*radius*radius. next you will need to know the formula for the circumfrence of a circle, which is 2*pi*r=Circ. You know that circumfrance is equal to 31.4 feet, so set that equal to the equation for circumfrence to find r. \[31.4 ft = 2*3.14*r\] because pi is equal to 3.14. Multiply 3.14 by 2 to get 6.28. 31.4 ft = 6.28*r divide 31.4 by 6.28 and you will get r. 5 ft = r. you will then take that information and put it into the equation to find the area: A=pi*r*r A=pi*5*5 can you figure out Area from there?

OpenStudy (anonymous):

no..I do not understand. Sorry

OpenStudy (anonymous):

thats fine, the only i can help you more is if you tell me if you dont understand. Ok so you are looking for area, and the formula for the area of a circle is always: Area= pi*radius*radius. but you are given circumfrance. do you have it up until there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, you are given the circumference of the circle, at 31.4 feet. The formula for the circumference of a circle is always C=2*pi*r. and you are given circumference, so set the equation to the value given to give you the formula: 31.4 feet = 2* pi * radius. Are we clear up until here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so we need to solve for radius because that is what we need for the equation of area. both 2 and pi are constants, so we can multiply pi by 2, and divide 31.4 feet by the product of 2 and pi, so the work will look like this: 31.4 feet = 2*pi*radius 31.4 feet = 6.28* radius (6.28 because has an accepted working value at 3.14) now divide 31.4 by 6.28 5= radius. Clear up until this point?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok, so you now know that the formula for area is A=pi*radius*radius. Solve for area now that you solved for radius. Can you figure this out?

OpenStudy (anonymous):

pi *5*5 = 3.14 x 5 x 5 = 81 ?

OpenStudy (anonymous):

not quite, you have the equation set up right, i would just punch that back into a calculator, because i am getting a number that is close to that, but not exactly

OpenStudy (anonymous):

Not using a calculator...lol

OpenStudy (anonymous):

78.5 my mistake...lol

OpenStudy (anonymous):

So that is my answer? 78.5?

OpenStudy (anonymous):

yes, the answer is 78.5

OpenStudy (anonymous):

and if I needed the diameter I would divide in half?

OpenStudy (anonymous):

nope, diameter is two times the radius. so you would multiply by 2

OpenStudy (anonymous):

but the formulas do not use diameter, so if you are given diameter, you must divide diameter by 2 to get the radius. dont use diameter when you have to use radius

OpenStudy (anonymous):

ok. Thanks. That's helped alot. Have to go to bed now. Big test tomorrow. Wish me luck :) Have a great night :)